# Asymptotically good CSS-T codes exist

### v1

@@ 2.1 Asymptotically good families of CSS-T codes
From any given asymptotically good family of binary CSS codes it is possible to derive an asymptotically good family of binary CSS-T codes.
@@
Examples of infinite asymptotically good families of binary CSS codes are presented, for instance, in [2, Theorem 1], [11, Theorem 1.2], and [28, Theorem 2]. The result follows from Theorem 4. ∎
2.2 Asymptotically good families of LDPC CSS-T codes
In this subsection, we show that the CSS-T construction of Theorem 3 preserves the Low-Density Parity Check (LDPC) property. An immediate application of this is the existence of asymptotically good quantum LDPC CSS-T codes. Quantum codes with sparse parity check matrices are especially interesting in the context of quantum computing. The LDPC property implies that the computation of each parity constraint imposed by the code involves a small number of qubits. As a result, the qubits of the code need to interact by means of a small number of two-qubit gates. Due to the fact that those gates used for syndrome extraction are also noisy, having a sparse amount of parity constraints is beneficial for practical implementations of quantum error correction codes [3, 5]. Furthermore, sparse parity check matrices are also important for decoding purposes [13, 34]. Thus, we discuss how our construction can be used to yield asymptotically good quantum LDPC codes. Before stating this result, we need some basic definitions and elementary results.
Let $C_{2}\subseteq C_{1}\subseteq\mathbb{F}_{2}^{n}$ be binary linear codes. Recall that the corresponding quantum CSS code has a parity check matrix
$H=\begin{bmatrix}H_{X}&0\\ 0&H_{Z}\end{bmatrix},$
where $H_{X}$ and $H_{Z}$ are generator matrices of $C_{2}$ and $C_{1}^{\perp}$, respectively. A quantum code with a sparse parity-check matrix $H$ is called a quantum LDPC code. Recall also that the following definition was given in Equation (2):
$C_{i}^{N_{I}}=\left\{(x,x):x\in C_{i}\right\}$
for each $i\in\{1,2\}$. In the following, we will use the shorter notation $C_{i}^{N}$ for $C_{i}^{N_{I}}$.
The next theorem shows that applying the result of Theorem 3 to a CSS code preserves its sparsity by sending its parity check matrix to another sparse parity check matrix.
@@ 2.2 Asymptotically good families of LDPC CSS-T codes
where $H_{X},H_{Z}$ generate $C_{2}$ and $C_{1}^{\perp}$, respectively. Then, the CSS-T code $\left(C_{1}^{N},C_{2}^{N}\right)$ has parity check matrix
@@
The definition of $C_{2}^{N}$ justifies that $H_{X}^{N}$ is a generator matrix for $C_{2}^{N}$. Moreover, one can easily check that the linear span of the rows of $H_{Z}^{N}$ is a subspace of $\left(C_{1}^{N}\right)^{\bot}$. Next, we show that the row space of $H_{Z}^{N}$ is in fact $(C_{1}^{N})^{\bot}$. Let $(x,y)\in\left(C_{1}^{N}\right)^{\perp}$ for some $x,y\in\mathbb{F}_{2}^{n}$. Let $e_{i}^{N}\vcentcolon=(e_{i},e_{i})$, where $e_{i}\in\mathbb{F}_{2}^{n}$ is the standard basis vector whose $j$-th entry is
$({e_{i}})_{j}=\begin{cases}1&\text{ if }i=j,\\ 0&\text{ if }i\neq j.\end{cases}$
So there exists $\alpha_{i}\in\mathbb{F}_{2}$ for each $1\leq i\leq n$ such that
@@
Now Lemma 2(3) implies that $x+y\in C_{1}^{\bot}$ and consequently $(x+y,0)$ is in the row space of $\begin{bmatrix}H_{Z}&0\end{bmatrix}$. Since $(y,y)$ belongs to the row space of $\begin{bmatrix}I&I\end{bmatrix}$, we conclude that $(x,y)$ is in the row space of $H_{Z}^{N}$.
The structure of $H_{X}^{N}$ shows that its maximum row weight is $2r_{X}$. Moreover, the last $n$ rows of $H_{Z}^{n}$ all have weight $2$. The maximum weight of the remaining rows of $H_{Z}^{N}$ is $r_{Z}$. Thus the maximum row weight of $H_{Z}^{N}$ is $\max\{r_{Z},2\}$. ∎
Recall that there exist several families of good quantum LDPC codes in the literature, for instance, see [14, 16, 23, 24, 29, 30]. An immediate consequence of the above theorem is that applying the construction of Theorem 3 to an LDPC CSS code preserves its sparsity. We use this property to entail the next result.
@@
A family of asymptotically good quantum LDPC code that is CSS-T exists.
@@
The proof follows by applying the quantum CSS-T construction of Theorem 3 to each asymptotically good quantum LDPC family presented above. Theorem 5 ensures that the resulting CSS-T codes are LDPC. ∎
3 Logical operators of different transversal gates
In this section, we first develop the idea used in the previous chapter to produce asymptotically good CSS codes for any given $Z$ rotation. Secondly, we show that for any $(C_{1},C_{2})_{\mathrm{CSS-T}}$ with an even code $C_{1}$, the logical operator corresponding to the transversal $T$ has multiplicative order four. Therefore, the transversal $T$ gate never realizes transversal logical $T$ gate on such CSS codes. Finally, we prove that the CSS-T codes from the construction of Theorem 3 also support transversal $CCZ$ gate. In this case, the corresponding logical operator is the identity.
3.1 Good CSS codes supporting transversal Z rotations
As we already stated in Corollary 1, one can construct a family of asymptotically good CSS codes that supports transversal $T$ gate. Obviously, the same family supports other larger $Z$ rotations, namely transversal $Z$ or $S$ gate. In this subsection, we discuss the construction of another asymptotically good CSS family that supports transversal application of other smaller $Z$ rotation gates.
First, recall that the $Z$ rotation gate by the angle $\frac{2\pi}{2^{\ell}}$ for some $\ell\geq 0$ is
Next, we show that applying the construction of Theorem 3 repeatedly preserves finer $Z$ rotations transversally.
@@ 3.1 Good CSS codes supporting transversal Z rotations
Let $u\in\mathbb{F}_{2}^{k}$, where $k$ is the number of logical qubits in the quantum code, and $H$ be a binary matrix having a basis of $C_{3}^{\ell}$ in its rows, where $C_{1}^{\ell}=C_{2}^{\ell}\oplus C_{3}^{\ell}$. Note that the number of physical qubits is $2^{\ell}n$. Next, we show that the following diagram commutes
|${\ket{u}_{L}}$${\displaystyle\sum_{v\in C_{2}^{l}}\ket{v+uH}}$|${\ket{u}_{L}}$${\displaystyle\sum_{v\in C_{2}^{l}}\varGamma^{(\ell)^{\otimes{2^{l}n}}}\ket{v+uH}}$$\scriptstyle{{\textrm{E}n}}$$\scriptstyle{I}$$\scriptstyle{\varGamma^{(\ell)^{\otimes{2^{\ell}n}}}}$$\scriptstyle{{\textrm{E}n}}$ (5)
The action of applying $\varGamma^{(\ell)}$ to all $2^{\ell}n$ physical qubits can be described as:
$\varGamma^{(\ell)^{\otimes{2^{\ell}n}}}\left(\sum_{v\in C_{2}^{\ell}}|v+uH\rangle\right)=\sum_{v\in C_{2}^{\ell}}\varGamma^{(\ell)^{\otimes{2^{\ell}n}}}|v+uH\rangle.$ (6)
Recall that the term $v+uH\in C_{1}^{\ell}$ for each selection of $u$ and $v$. Now, applying the operator $\varGamma^{(\ell)^{\otimes{2^{\ell}n}}}$ to each term in Equation (6) gives
$\sum_{v\in C_{2}^{\ell}}\varGamma^{(\ell)^{\otimes{2^{\ell}n}}}|v+uH\rangle=\sum_{v\in C_{2}^{\ell}}\left(\mathrm{e}^{\frac{2\pi i}{2^{\ell}}}\right)^{{\textrm{w}t}{(v+uH)}}|v+uH\rangle,$
where ${\textrm{w}t}{(v+uH)}$ is the weight of corresponding vector in $C_{1}^{\ell}$. Moreover, since $C_{1}^{\ell}$ is obtained from applying the procedure of Theorem 3 $\ell$ times, its codewords are in the form
$C_{1}^{\ell}=\{(x,x,\ldots,x):x\in C_{1}\},$
where the number of repetitions is $2^{\ell}$ times. Hence the weight of each codeword of $C_{1}^{\ell}$ is divisible by $2^{\ell}$. Thus
$\sum_{v\in C_{2}^{\ell}}\varGamma^{(\ell)^{\otimes{2^{\ell}n}}}|v+uH\rangle=\sum_{v\in C_{2}^{\ell}}\left(\mathrm{e}^{\frac{2\pi i}{2^{l}}}\right)^{{\textrm{w}t}{(v+uH)}}|v+uH\rangle=\sum_{v\in C_{2}^{\ell}}|v+uH\rangle=|u\rangle_{L}.$
This shows that the diagram of Equation (5) commutes, and completes the proof. ∎
Note that one may get the same result, with a different logical operator, after applying Theorem 3 to $C_{2}\subseteq C_{1}$ less than $\ell$ times. However, using $\ell$ repetitions keeps the proof more intuitive and simpler. An application of Theorem 6 is the following interesting result. The proof is similar to that of Corollary 1. So we skip it in here.
@@
Let $\ell>0$ be a positive integer. Then a family of asymptotically good CSS code exists that supports transversal $\varGamma^{(\ell)}$ gate.
3.2 Logical operator of CSS-T codes with even weights
Recall that applying the $T$ gate transversally to all the physical qubits of a CSS-T code realizes a logical gate. In this subsection, we investigate the structure of this logical operator when the $C_{2}\subseteq C_{1}$ both are even and form a CSS-T code. We show that the multiplicative order of such a logical operator is always a divisor of four. In particular, such CSS-T codes never realize a transversal logical $T$ gate as it has the multiplicative order of eight.
@@ 3.2 Logical operator of CSS-T codes with even weights
Recall that $C_{2}^{N}\subseteq C_{1}^{N}$ is the pair of classical codes used in Theorem 3. Their definition shows that $C_{1}^{N}$ also has only even weight codewords. Therefore, the logical operator corresponding to the transversal $T$, when applied to the quantum code of Theorem 3, has order a factor of four. All the CSS-T codes obtained from the extended construction of classical codes and the CSS-T Reed–Muller (RM) (but not punctured RM) codes satisfy the conditions of Theorem 7.
3.3 Application of transversal CCZ gate
@@
In this subsection, we show that the quantum CSS code obtained from Theorem 3 also support transversal CCZ gate. First, recall that for each $C_{2}\subseteq C_{1}$, we have set
$C_{i}^{N}\vcentcolon=\{(z,z):z\in C_{i}\}$
for each $i\in\{1,2\}$. We also fix $H$ to be a binary matrix having a basis of $C_{3}^{N}$ in its rows, where $C_{1}^{N}=C_{2}^{N}\oplus C_{3}^{N}$.
@@
Let $C_{2}\subseteq C_{1}\subseteq\mathbb{F}_{2}^{n}$ be two binary codes. The quantum CSS-T code $\left(C_{1}^{N},C_{2}^{N}\right)_{\mathrm{CSS-T}}$ has the following property: applying the transversal $CCZ$ gate to all $2n$ physical qubits realizes the logical identity gate.
@@
$CCZ^{\otimes 2n}\left(\sum_{v\in C_{2}^{N}}|v+uH\rangle\right)\left(\sum_{v^{\prime}\in C_{2}^{N}}|v^{\prime}+xH\rangle\right)\left(\sum_{v^{\prime\prime}\in C_{2}^{N}}|v^{\prime\prime}+wH\rangle\right).$ (9)
@@
$\begin{split}&\sum_{v,v^{\prime},v^{\prime\prime}\in C_{2}^{N}}CCZ^{\otimes 2n}|v+uH\rangle|v^{\prime}+xH\rangle|v^{\prime\prime}+wH\rangle\\ &=\sum_{v,v^{\prime},v^{\prime\prime}\in C_{2}^{N}}CCZ^{\otimes 2n}|a\rangle|b\rangle|c\rangle=\sum_{v,v^{\prime},v^{\prime\prime}\in C_{2}^{N}}(-1)^{\displaystyle\sum_{i=1}^{2n}a_{i}b_{i}c_{i}}|a\rangle|b\rangle|c\rangle,\end{split}$
where $a=v+uH=(\alpha,\alpha)$, $b=v^{\prime}+xH=(\beta,\beta)$, and $c=v^{\prime\prime}+wH=(\gamma,\gamma)$ are codewords of $C_{1}^{N}$ for some $\alpha,\beta,\gamma\in\mathbb{F}_{2}^{n}$. Thus
$\sum_{i=1}^{2n}a_{i}b_{i}c_{i}=2\sum_{i=1}^{n}\alpha_{i}\beta_{i}\gamma_{i}\equiv 0\pmod{2}.$
This implies that $(-1)^{\displaystyle\sum_{i=1}^{2n}a_{i}b_{i}c_{i}}=1$. Thus, the result of (9) is
@@
This implies that $CCZ^{\otimes 2n}$ acts as identity on physical qubits. This completes the proof. ∎

### v2

@@ III-A Asymptotically good families of CSS-T codes
From any given asymptotically good family of binary CSS codes, it is possible to derive an asymptotically good family of binary CSS-T codes.
@@
Infinite asymptotically good families of binary CSS codes are presented in e.g., [26, Theorem 1], [27, Theorem 1.2], and [28, Theorem 2]. The result then follows from Theorem 5. ∎
III-B Asymptotically good families of LDPC CSS-T codes
Quantum codes with low-density parity check (LDPC) matrices are especially interesting in the context of quantum computing. Here we discuss how Theorem 4 can yield asymptotically good quantum LDPC codes. The next theorem shows that applying the result of Theorem 4 to a CSS code preserves its sparsity by sending its parity check matrix to another sparse parity check matrix.
@@ III-B Asymptotically good families of LDPC CSS-T codes
where $H_{X},H_{Z}$ generate $C_{2}$ and $C_{1}^{\perp}$, respectively. Then, $\left(C_{1}^{N},C_{2}^{N}\right)_{\text{CSS-T}}$ has parity check matrix
@@
The definition of $C_{2}^{N}$ justifies that $H_{X}^{N}$ is a generator matrix for $C_{2}^{N}$. Moreover, one can easily check that the linear span of the rows of $H_{Z}^{N}$ is a subspace of $\left(C_{1}^{N}\right)^{\bot}$. Next, we show that the row space of $H_{Z}^{N}$ is $(C_{1}^{N})^{\bot}$. Let $(x,y)\in\left(C_{1}^{N}\right)^{\perp}$ for some $x,y\in\mathbb{F}_{2}^{n}$. Let $e_{i}^{N}\vcentcolon=(e_{i},e_{i})$, where $e_{i}\in\mathbb{F}_{2}^{n}$ is the standard basis vector.So there exists $\alpha_{i}\in\mathbb{F}_{2}$ for each $1\leq i\leq n$ such that
@@
This implies that $\forall a^{N}=(a,a)\in C_{1}^{N},(x+y,0)\cdot a^{N}=0$, hence $\forall a\in C_{1},(x+y)\cdot a=0$, and thus $x+y\in C_{1}^{\bot}$. Consequently, $(x+y,0)$ is in the row space of $\begin{bmatrix}H_{Z}&0\end{bmatrix}$. Since $(y,y)$ belongs to the row space of $\begin{bmatrix}I&I\end{bmatrix}$, $(x,y)$ is in the row space of $H_{Z}^{N}$. The structure of $H_{X}^{N}$ shows that its maximum row weight is $2r_{X}$. Moreover, the last $n$ rows of $H_{Z}^{n}$ have all weight $2$. The maximum weight of the remaining rows of $H_{Z}^{N}$ is $r_{Z}$. Thus, the maximum row weight of $H_{Z}^{N}$ is $\max\{r_{Z},2\}$. ∎
There exist several families of good quantum LDPC codes, see e.g., [29, 30, 31, 32, 33, 34]. An immediate consequence of the above theorem is that applying the construction of Theorem 4 to an LDPC CSS code preserves its sparsity. We use this property to entail the next result.
@@
A family of asymptotically good quantum LDPC codes that are CSS-T exists.
@@
We apply the CSS-T construction of Theorem 4 to each asymptotically good quantum LDPC family presented above. Theorem 6 ensures that the resulting CSS-T codes are LDPC. ∎
IV Logical operators of different transversal gates
In this section, we first develop the idea used in the previous chapter to produce asymptotically good CSS codes for any given $Z$ rotation. Then we discuss an application of them in mitigating coherent noise. Secondly, we show that for any $(C_{1},C_{2})_{\mathrm{CSS-T}}$ with an even code $C_{1}$, the logical operator corresponding to the transversal $T$ has multiplicative order four. Therefore, the transversal $T$ gate never realizes transversal logical $T$ gate on such CSS codes. Finally, we prove that the CSS-T codes from the construction of Theorem 4 also support transversal $CCZ$ gate. In this case, the corresponding logical operator is the identity.
IV-A Good CSS codes supporting transversal Z rotations
As we already stated in Corollary 1, one can construct a family of asymptotically good CSS codes that supports transversal $T$ gate. Obviously, the same family supports other larger $Z$ rotations, namely transversal $Z$ or $S$ gate. In this subsection, we discuss the construction of another asymptotically good CSS family that supports transversal application of other smaller $Z$ rotation gates. Next, we show that applying the construction of Theorem 4 repeatedly preserves finer $Z$ rotations transversally.
@@ IV-A Good CSS codes supporting transversal Z rotations
Let $u\in\mathbb{F}_{2}^{k}$, where $k$ is the number of logical qubits in the quantum code, and $H$ be a binary matrix having a basis of $C_{3}^{\ell}$ in its rows, where $C_{1}^{\ell}=C_{2}^{\ell}\oplus C_{3}^{\ell}$. Note that the number of physical qubits is $2^{\ell}n$. Consider the logical state |$\ket{u}_{L}=\sum_{v\in C_{2}^{l}}\ket{v+uH}$. Recall that the term $v+uH\in C_{1}^{\ell}$ for each selection of $u$ and $v$ and ${\textrm{w}t}{(v+uH)}$ is the weight of corresponding vector in $C_{1}^{\ell}$. Moreover, since $C_{1}^{\ell}$ is obtained from applying the procedure of Theorem 4 $\ell$ times, its codewords have weights divisible by $2^{\ell}$. Thus
$\begin{split}\varGamma^{(\ell)^{\otimes{2^{\ell}n}}}\left(\sum_{v\in C_{2}^{\ell}}|v+uH\rangle\right)&=\sum_{v\in C_{2}^{\ell}}\varGamma^{(\ell)^{\otimes{2^{\ell}n}}}|v+uH\rangle\\ &=\sum_{v\in C_{2}^{\ell}}\left(\mathrm{e}^{\frac{2\pi i}{2^{\ell}}}\right)^{{\textrm{w}t}{(v+uH)}}|v+uH\rangle\\ &=\ket{u}_{L}.\end{split}$ (3)
The proof of the previous Theorem 7 can partially follow from [35, Remark 13]. However, the latter does not provide any information about the structure of the logical operator. Note that one may get the same result after applying Theorem 4 to $C_{2}\subseteq C_{1}$ less than $\ell$ times. However, using $\ell$ repetitions keeps the proof more intuitive. An application of Theorem 7 is the following interesting result. The proof is similar to that of Corollary 1. So we skip it in here.
@@
Let $\ell>0$ be a positive integer. Then, a family of asymptotically good CSS code exists that supports transversal $\varGamma^{(\ell)}$ gate.
IV-A1 Asymptotically good CSS codes for mitigating restricted coherent noise
In quantum systems, coherent errors refer to systematic rotations around a fixed axis, which can accumulate over time, potentially leading to more severe damage than stochastic errors. A natural and commonly studied model considers coherent noise acting transversally, where each qubit is subjected to an independent unitary operation [1]. This is the case, for example, when a uniform background magnetic field with the Hamiltonian $H=Z_{1}+Z_{2}+\cdots+Z_{n},$ acts on the system, where $Z_{i}$ denotes the Pauli-$Z$ operator acting on the $i$-th qubit. Such interaction can induce a coherent error, corresponding to a $Z$-rotation by an angle $\frac{2\pi}{2^{\ell}}$, for some integer $\ell>0$, on each qubit. Although coherent noise can be addressed using active error correction techniques, a more resource-efficient approach is to passively mitigate its effects by encoding logical information into decoherence-free subspaces (DFSs) [36, 37]. Particularly, in the language of stabilizer codes, the noise is required to preserve the code space and act as the logical identity.
The structure of CSS codes for which each arbitrary transversal $Z$-rotations act as the logical identity is discussed in [1], and a construction for such codes is provided. These codes are commonly referred to as CSS codes that are oblivious to coherent noise.
Let $\ell_{\max}\geq 3$ be an integer. In this paper, we consider CSS codes where the code space is preserved by transversal $\frac{2\pi}{2^{\ell}}$ $Z$-rotation, which acts as the identity operator on the code space, for each $\ell\leq\ell_{\max}$. We refer to such codes as oblivious to $\frac{2\pi}{2^{\ell_{\max}}}$-coherent noise. This approach is in line with the research direction presented in [1], and in the following theorem we give a construction of asymptotically good quantum CSS codes that are oblivious to $\frac{2\pi}{2^{\ell_{\max}}}$-coherent noise, for any $\ell_{\max}$.
Theorem 8.
Let $\ell_{max}\geq 3$ be an integer. Then the binary quantum CSS code $\left(C_{1}^{\ell_{\max}},C_{2}^{\ell_{\max}}\right)$, obtained from applying Theorem 4 $\ell_{\max}$-times to $C_{2}\subseteq C_{1}\subseteq\mathbb{F}_{2}^{n}$ is oblivious to $\frac{2\pi}{2^{\ell_{\max}}}$-coherent noise.
Proof.
The proof of Theorem 7 shows that transversal $\frac{2\pi}{2^{\ell_{\max}}}$ $Z$-rotation implies the logical identity operator on the encoded space. Applying this transversal gate $2^{\ell_{m}-\ell}$ times, for each $\ell\leq\ell_{\max}$, shows that again the corresponding logical operator is the identity, which is in correspondence to transversal $\frac{2\pi}{2^{\ell}}$ $Z$-rotation. Hence such CSS code is oblivious to $\frac{2\pi}{2^{\ell_{\max}}}$-coherent noise. $\frac{2\pi}{2^{\ell_{\max}}}$-coherent noise. ∎
IV-B Logical operator of CSS-T codes with even weights
Recall that applying the $T$ gate transversally to all the physical qubits of a CSS-T code realizes a logical gate. In this subsection, we investigate the structure of this logical operator when both $C_{2}\subseteq C_{1}$ are even and form a CSS-T code. We show that the multiplicative order of such a logical operator is always a divisor of four. In particular, such CSS-T codes never realize a transversal logical $T$ gate as the latter has the multiplicative order of eight.
@@ IV-B Logical operator of CSS-T codes with even weights
Thus, all CSS-T codes obtained from the extended construction of classical codes and Reed–Muller (RM) (but not punctured RM) codes satisfy the conditions of Theorem 9.
IV-C Application of transversal CCZ gate
@@
In this subsection, we show that the quantum CSS code obtained from Theorem 4 also supports the transversal CCZ gate. Recall that for each $C_{2}\subseteq C_{1}$, we have set $C_{i}^{N}\vcentcolon=\{(z,z):z\in C_{i}\}$ for $i\in\{1,2\}$. We also fix $H$ to be a binary matrix having a basis of $C_{3}^{N}$ in its rows, where $C_{1}^{N}=C_{2}^{N}\oplus C_{3}^{N}$.
@@
Let $C_{2}\subseteq C_{1}\subseteq\mathbb{F}_{2}^{n}$ be two binary codes. The quantum CSS-T code $\left(C_{1}^{N},C_{2}^{N}\right)_{\mathrm{CSS}}$ has the following property: applying the transversal $CCZ$ gate to all $2n$ physical qubits realizes the logical identity gate.
@@
$\begin{split}CCZ^{\otimes 2n}&\left(\sum_{v\in C_{2}^{N}}|v+uH\rangle\right)\left(\sum_{v^{\prime}\in C_{2}^{N}}|v^{\prime}+xH\rangle\right)\\ &\left(\sum_{v^{\prime\prime}\in C_{2}^{N}}|v^{\prime\prime}+wH\rangle\right).\end{split}$ (6)
@@
$\begin{split}&\sum_{v,v^{\prime},v^{\prime\prime}\in C_{2}^{N}}CCZ^{\otimes 2n}|v+uH\rangle|v^{\prime}+xH\rangle|v^{\prime\prime}+wH\rangle\\ &=\sum_{v,v^{\prime},v^{\prime\prime}\in C_{2}^{N}}CCZ^{\otimes 2n}|a\rangle|b\rangle|c\rangle\\ &=\sum_{v,v^{\prime},v^{\prime\prime}\in C_{2}^{N}}(-1)^{\displaystyle\sum_{i=1}^{2n}a_{i}b_{i}c_{i}}|a\rangle|b\rangle|c\rangle\\ &=\sum_{v,v^{\prime},v^{\prime\prime}\in C_{2}^{N}}(-1)^{\displaystyle 2\sum_{i=1}^{n}\alpha_{i}\beta_{i}\gamma_{i}}|a\rangle|b\rangle|c\rangle,\end{split}$
where $a=v+uH=(\alpha,\alpha)$, $b=v^{\prime}+xH=(\beta,\beta)$, and $c=v^{\prime\prime}+wH=(\gamma,\gamma)$ are codewords of $C_{1}^{N}$ for some $\alpha,\beta,\gamma\in\mathbb{F}_{2}^{n}$. Thus, the result of (6) is
@@
This implies that $CCZ^{\otimes 2n}$ acts as the logical identity. ∎